Use the facts that log 5 = 0.6990 and log 7 = 0.8451 to find the value of the logarithm. log 35
Answer by htmentor(1343)

Use the facts that log 5 = 0.6990 and log 7 = 0.8451 to find the value of the logarithm. log 35
=====================
Using the proprety log(a*b) = log(a) + log(b), we have
log(35) = log(5*7) = log(5) + log(7) = 0.699 + 0.8451 = 1.5441